OpenStudy (anonymous):

Find the solution of the differential equation y'= 2x/(1+x^4) , such that y(0)=0

OpenStudy (mimi_x3):

\[\frac{dy}{dx}=\frac{2x}{1+x^4}\] just integrate both sides i suppose..

OpenStudy (mimi_x3):

\[\int \frac{dy}{dx} = \int \frac{2x}{1+x^4}\]

OpenStudy (anonymous):

if \[\Large u= 1 + x^4 \] then \[\Large \frac{du}{dx}=4x^3 \] How would you continue there? I am just curious, please don't misunderstand this, I know that there are plenty of ways to solve such an equation.

OpenStudy (mimi_x3):

I made a mistake.. trig substitution..or \(u=x^2\)

OpenStudy (zugzwang):

\[\frac{dy}{dx}=\frac{2x}{1+x^4}\]\[dy=\frac{2x}{1+x^4}dx\]

OpenStudy (zugzwang):

\[\int\limits_{}^{} dy=\int\limits_{}^{}\frac{2x}{1+x^4}dx\]